# Decoding the string in terms of holographic quantum maps

### v1

@@ Abstract
It has recently been shown that the Nambu-Goto equation for a string emerges from the junction conditions in three-dimensional gravity. Holographically, gravitational junctions are dual to interfaces in conformal field theory. We demonstrate that each stringy mode of the junction corresponds to a universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum map between in and out Hilbert spaces of excitations scattered at the interface, and also a universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum map relating the excitations on both sides. These quantum maps generalize those realized by defect operators and preserve the conformal boundary condition at the interface.
Introduction:- The holographic duality states that quantum gravity, particularly a superstring theory, can be reformulated in terms of a nongravitational quantum field theory living at the boundary of spacetime [1, 2, 3]. Currently, there has been a lot of progress in our understanding of how the bulk spacetime and semi-classical bulk effective theory consisting of Einstein’s gravity coupled to a few fields can be reconstructed from the boundary quantum field theory in the large N and strong coupling limit [4, 5, 6, 7]. The reconstruction of dynamical extended objects (branes) of the gravitational theory that are necessary for its non-perturbative completion, is therefore a problem of fundamental importance.
@@
Recently, it has been shown that the Nambu-Goto equation for the string emerges from the gravitational junction conditions in three dimensions [19]. In the present letter, we report how these stringy modes of gravitational junctions correspond to universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum maps from the in to out Hilbert spaces of excitations scattered at the dual interface. Equivalently, the correspondence can be formulated via universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum maps relating the Hilbert spaces of the two CFTs that straddle the interface. These maps generalize the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ and $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ maps realized by a defect operator while preserving the conformal boundary condition at the interface. We also prove the universality of the quantum maps corresponding to the holographic defect operator.
The conformal interface:- The non-vanishing components of the energy-momentum tensor $T_{\mu\nu}$ in any state of a two-dimnsional CFT are $T_{\pm\pm}=T_{\mu\nu}n_{\pm}^{\mu}n_{\pm}^{\nu}$, with $n_{\pm}^{\mu}$ being the two future directed null vectors. The integration of $T_{++}$ and $T_{--}$ at past null infinity define the operators $\mathcal{E}_{\pm}$ which measure the left-moving and right-moving incoming energy fluxes, respectively. A conformal interface is a gluing of two CFTs, namely CFTL and CFTR on the left and right, respectively, such that scale invariance is preserved. The interface at $x=0$ is represented by an operator insertion $I_{L,R}$ which satisfies
$(L_{n,+}^{L}-L_{-n,-}^{L})I_{L,R}=I_{L,R}(L_{n,+}^{R}-L_{-n,-}^{R})$ (1)
@@
It is possible to generalize the conformal interface with a half-sided conformal transformation in which the same conformal transformation is applied to both the left and right movers on the left or the right side. Such a transformation preserves the interface at $x=0$ and the $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ map is then simply composed with a conformal transformation on CFTR. As a result, the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map is redefined. We will discuss these explicitly later.
@@
Furthermore, it has been shown in [19], when $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are copies of an Einstein manifold $\mathcal{M}$, that the general solutions of the junction conditions are in one-to-one correspondence with solutions of the non-linear Nambu-Goto equation for a worldsheet in $\mathcal{M}$. Specifically, the hypersurface
@@
$16\sigma\partial^{2}_{\tau}x_{s}+4(4-\lambda^{2})\left(2\partial_{\sigma}x_{s}-\sigma\partial_{\sigma}^{2}x_{s}\right)=-ie^{i\omega\tau}(-4+\lambda^{2})\sigma^{3}\omega\left((L_{-}^{R}-L_{+}^{L})e^{\frac{i\lambda\sigma}{\sqrt{4-\lambda^{2}}}}+(L_{-}^{L}-L_{+}^{R})e^{-\frac{i\lambda\sigma}{\sqrt{4-\lambda^{2}}}}\right),$ (14)
@@
$L_{+}^{L}=\mathcal{T}_{L}L_{+}^{R}+(1-\mathcal{T}_{R})L_{-}^{L},$ (19)
$L_{-}^{R}=\mathcal{T}_{R}L_{-}^{L}+(1-\mathcal{T}_{L})L_{+}^{R},$
@@
where $\alpha$ is an arbitrary constant parameter which acts as an overall (diagonal) conformal transformation that changes the background state (see below). It is easy to see using (24) that (Decoding the string in terms of holographic quantum maps) preserves the interface at $\tilde{x}_{1,2}=0$ where $\tilde{t}_{2}=\tilde{t}_{1}$. We therefore obtain continuous coordinates and metric across the interface. These conformal transformations can be uplifted to bulk diffeomorphisms ($X^{\mu}\rightarrow\widetilde{X}^{\mu}$ with $X^{\mu}$ denoting bulk coordinates) on both sides [28, 26, 30], and both the induced metric and the extrinsic curvatures of $\Sigma_{1,2}$ remain invariant under these bulk diffeomorphisms ($X^{\mu}(\tau,\sigma)\rightarrow\widetilde{X}^{\mu}(\tau,\sigma)$), producing an equivalent solution of the junction conditions.
@@
$\quad-\frac{L_{-}^{L}}{4}(\alpha)\left(-2+\mathcal{T}_{R}(2+\lambda)\right)+\frac{\alpha a_{n}}{2},$ (28)
@@
These transformations are reproduced in the holographic renormalization procedure of extracting the dual energy-momentum tensors via the bulk diffeomorphisms which uplift the corresponding conformal transformations [30]. The transformed amplitudes also satisfy the energy conservation (and thus also conformal boundary condition) at the linear order as
@@
$g_{+}(x^{+})=x^{+}-\epsilon\frac{e^{i\omega x^{+}}}{\omega^{3}}\frac{4ia_{n}}{2-\lambda}+\mathcal{O}(\epsilon^{2}),$
$g_{-}(x^{-})=x^{-}+\epsilon\frac{e^{i\omega x^{-}}}{\omega^{3}}\frac{4ia_{n}}{2-\lambda}+\mathcal{O}(\epsilon^{2}).$ (32)
on the left and right movers, respectively. It also follows from (26), (27), (28) and (29) that the physical fluxes satisfy
@@
so that the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map can be rewritten in the form $\widetilde{\mathcal{D}}\circ\mathcal{S}$ where $\widetilde{\mathcal{D}}$ is a redistribution of energy among the two outgoing energy fluxes which can be realized by a linearized conformal transformation like $\mathcal{D}$. The $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map in both forms, namely $\mathcal{S}\circ\mathcal{D}$ and $\widetilde{\mathcal{D}}\circ\mathcal{S}$, is independent of the background state (represented by the choice of $\alpha$) and reduces to the universal scattering matrix $S$ given by (3) in absence of the stringy mode $a_{n}$, which otherwise gives rise to universal energy redistribution among the incoming or outgoing energy fluxes.
We also note from (26), (27), (28) and (29) that the physical fluxes satisfy
@@
We also note that for any value of $\lambda$ between 0 and $2$, the interface is topological when $a_{n}$ satisfies
$a_{n}=\frac{\lambda}{2}\left(\tilde{L}_{-}^{L}-\tilde{L}_{+}^{R}\right)$ (36)
as in this case $\tilde{L}_{\pm}^{R}=\tilde{L}_{\pm}^{L}$.
Finally, the achronal ANEC is always satisfied for the generalized quantum maps at the linear order as the linearized conformal transformations do not change the energy fluxes if $a_{n}(\omega)$ vanishes as $\omega\rightarrow 0$ (see (Decoding the string in terms of holographic quantum maps) as for instance). [^1] Under $g(x^{\pm})=x^{\pm}+\epsilon\mathbb{g}(x^{\pm})$, $\int{\rm d}x^{\pm}T_{\pm\pm}$ is invariant at $\mathcal{O}(\epsilon)$ if $\int{\rm d}x^{\pm}\mathbb{g}^{\prime\prime\prime}(x^{+})=0$ when $T_{\pm\pm}$ are also $\mathcal{O}(\epsilon)$. For $\mathbb{t}_{d}$ and other variables to be well defined, we actually require that $a_{n}(\omega)$ vanishes at least as $\omega^{3}$ as $\omega\rightarrow 0$.
@@
Conclusions:- In this letter, we have demonstrated that the gravitational two-way junction including its stringy degrees of freedom can be decoded in terms of quantum maps. In absence of non-trivial stringy modes, these maps reduce to frequency independent universal maps [20, 31, 32, 21] that are realized by a defect operator, while the stringy modes imply that these maps should be generalized by composing them with appropriate conformal transformations on left and right movers. The full quantum maps preserve the conformal boundary condition on the interface. Particularly, the energy scattering involves independent and equal transmissions/reflections from both sides composed with an energy redistribution among the incoming/outgoing fluxes realized by conformal transformations. We have demonstrated the universality of these quantum maps by showing that they work for arbitrary input fluxes and an arbitrary background state (instead of the vacuum) on both sides of the interface within the universal sector.
@@
The understanding of the full non-linear gravitational problem, involving mixing of modes with different frequencies, in terms of quantum maps will be of fundamental importance for a deeper understanding of holographic reconstruction. It would be also of profound interest to understand how explicit reconstruction of sub-regions of the gravitational junction from interface CFT sub-regions can work by generalizing the reformulation of bulk reconstruction in holography in terms of recovery maps of quantum error correcting codes [4, 7]. As a first step, it would be of interest to understand how stringy degrees of freedom of gravitational junctions influence entanglement in the dual CFT following [33, 34, 35].
@@
It would also be interesting to investigate the quantum null energy condition (QNEC) [36, 37, 38] in the holographic interfaces dual to gravitational junctions. Recently, it has been shown that the QNEC can impose non-trivial quantum thermodynamic bounds even if matter localized on bulk hypersurfaces satisfies the classical null energy conditions [39, 40, 41] (see also [42] for a similar study in CFTs). The understanding of the link between quantum thermodynamic bounds and energy conditions is fundamental for the realization of quantum engines and processors using interfaces in many-body systems.
@@ Acknowledgements.
Acknowledgments:- We thank Costas Bachas, Avik Banerjee and Giuseppe Policastro for valuable discussions. We especially thank Giuseppe Policastro for discussions related to the ANEC bounds and quantum maps. AC, AM and MM acknowledge support from FONDECYT postdoctoral grant no. 3230222, FONDECYT regular grant no. 1240955 and “Doctorado Nacional” grant no. 21250596 of La Agencia Nacional de Investigación y Desarrollo (ANID), Chile, respectively. AM expresses appreciation for the hospitality of LPENS Paris where a major portion of this work has been carried out.

### v2

@@ Abstract
It has recently been shown that the Nambu-Goto equation for a string emerges from the junction conditions in three-dimensional gravity. Holographically, gravitational junctions are dual to interfaces in conformal field theory. We demonstrate that each stringy mode of the junction corresponds to a universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum map between in and out Hilbert spaces of excitations scattered at the interface, and also a universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum map relating the excitations on both sides. These quantum maps generalize those realized by defect operators, preserve the conformal boundary condition at the interface, and realize a tunable energy transmitter.
Introduction:- The holographic duality states that quantum gravity, particularly a superstring theory, can be reformulated in terms of a nongravitational quantum field theory living at the boundary of spacetime [1, 2, 3]. There has been a lot of progress in our understanding of how the bulk spacetime and semi-classical bulk effective theory consisting of Einstein’s gravity coupled to a few fields can be reconstructed from the boundary quantum field theory in the large N and strong coupling limit [4, 5, 6, 7]. The reconstruction of dynamical extended objects (branes) of the gravitational theory that are necessary for its non-perturbative completion, is therefore a problem of fundamental importance.
@@
It has been shown that the Nambu-Goto equation for the string emerges from the gravitational junction conditions in three dimensions [19]. In this letter, we report how these stringy modes of gravitational junctions correspond to universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum maps from the in to out Hilbert spaces of excitations scattered at the dual interface. Equivalently, the correspondence can be formulated via universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum maps relating the Hilbert spaces of the two CFTs that straddle the interface. These maps generalize the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ and $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ maps realized by a defect operator while preserving the conformal boundary condition at the interface. We also prove the universality of the quantum maps corresponding to the holographic defect operator.
The conformal interface:- The non-vanishing components of the energy-momentum tensor $T_{\mu\nu}$ in any state of a two-dimensional CFT are $T_{\pm\pm}=T_{\mu\nu}n_{\pm}^{\mu}n_{\pm}^{\nu}$, with $n_{\pm}^{\mu}$ being the two future directed null vectors. The integration of $T_{++}$ and $T_{--}$ at past null infinity define the operators $\mathcal{E}_{\pm}$ which measure the left-moving and right-moving incoming energy fluxes, respectively. A conformal interface is a gluing of two CFTs, namely CFTL and CFTR on the left and right, respectively, such that scale invariance is preserved. The interface at $x=0$ is represented by an operator insertion $I_{L,R}$ which satisfies
$(L_{n,+}^{L}-L_{-n,-}^{L})I_{L,R}=I_{L,R}(L_{n,+}^{R}-L_{-n,-}^{R}),$ (1)
@@
It is possible to generalize the conformal interface with a half-sided conformal transformation in which the same conformal transformation is applied to the left and right movers on either the left or the right side. Such a transformation preserves the interface at $x=0$ and the $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ map is then simply composed with a conformal transformation on CFTR, which results in a redefinition of the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map. We will discuss these explicitly later.
@@
It has been shown in [19], when $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are copies of an Einstein manifold $\mathcal{M}$, that the general solutions of the junction conditions are in one-to-one correspondence with solutions of the non-linear Nambu-Goto equation for a worldsheet in $\mathcal{M}$. Specifically, the hypersurface
@@
$16\sigma\partial^{2}_{\tau}x_{s}+4(4-\lambda^{2})\left(2\partial_{\sigma}x_{s}-\sigma\partial_{\sigma}^{2}x_{s}\right)=-ie^{i\omega\tau}(-4+\lambda^{2})\sigma^{3}\omega\left((L_{-}^{R}-L_{+}^{L})e^{\frac{i\lambda\omega\sigma}{\sqrt{4-\lambda^{2}}}}+(L_{-}^{L}-L_{+}^{R})e^{-\frac{i\lambda\omega\sigma}{\sqrt{4-\lambda^{2}}}}\right),$ (14)
@@
$L_{+}^{L}=\mathcal{T}_{L}L_{+}^{R}+(1-\mathcal{T}_{R})L_{-}^{L},\quad L_{-}^{R}=\mathcal{T}_{R}L_{-}^{L}+(1-\mathcal{T}_{L})L_{+}^{R},$ (19)
@@
where $\alpha$ is an arbitrary constant parameter which acts as an overall (diagonal) conformal transformation that changes the background state (see below). It is easy to see using (24) that (Decoding the string in terms of holographic quantum maps) preserves the interface at $\tilde{x}_{1,2}=0$ where $\tilde{t}_{2}=\tilde{t}_{1}$. We therefore obtain continuous coordinates and metric across the interface. These conformal transformations can be uplifted to bulk diffeomorphisms ($X^{\mu}(\tau,\sigma)\rightarrow\widetilde{X}^{\mu}(\tau,\sigma)$ with $X^{\mu}$ denoting bulk coordinates on both sides [28, 26, 30], and both the induced metric and the extrinsic curvatures of $\Sigma_{1,2}$ remain invariant under these bulk diffeomorphisms, producing an equivalent solution of the junction conditions.
@@
$\quad-\frac{L_{-}^{L}}{4}\alpha\left(-2+\mathcal{T}_{R}(2+\lambda)\right)+\frac{\alpha a_{n}}{2},$ (28)
@@
These transformations are reproduced in the holographic renormalization procedure of extracting the dual energy-momentum tensors via the bulk diffeomorphisms which uplift the corresponding conformal transformations [30]. The transformed amplitudes also satisfy energy conservation (and thus also conformal boundary condition) at linear order as
@@
$g_{\pm}(x^{\pm})=x^{\pm}\mp\epsilon\frac{e^{i\omega x^{\pm}}}{\omega^{3}}\frac{4ia_{n}}{2-\lambda}+\mathcal{O}(\epsilon^{2}).$ (32)
The physical fluxes (26), (27), (28) and (29) also satisfy
@@
so that the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map can be rewritten in the form $\widetilde{\mathcal{D}}\circ\mathcal{S}$ where $\widetilde{\mathcal{D}}$ is a redistribution of energy among the two outgoing energy fluxes which can also be realized by a linearized conformal transformation. The $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map in both forms, namely $\mathcal{S}\circ\mathcal{D}$ and $\widetilde{\mathcal{D}}\circ\mathcal{S}$, is independent of the background state (represented by the choice of $\alpha$) and reduces to the universal scattering matrix $S$ given by (3) in absence of the stringy mode $a_{n}$, which otherwise gives rise to universal energy redistribution among the incoming or outgoing energy fluxes.
We also note from (26), (27), (28) and (29) that
@@
For any value of $\lambda$ between 0 and $2$, the interface is topological when
$a_{n}=\frac{\lambda}{2}\left(\tilde{L}_{-}^{L}-\tilde{L}_{+}^{R}\right).$ (36)
In this case, as evident from (Decoding the string in terms of holographic quantum maps), $\tilde{L}_{\pm}^{R}=\tilde{L}_{\pm}^{L}$ so that we obtain perfect energy transmission. Similarly, the interface is always factorizing when
$a_{n}=\tilde{L}_{+}^{R}-\tilde{L}_{-}^{L}$ (37)
as in this case $\tilde{L}_{+}^{R,L}=\tilde{L}_{-}^{R,L}$ and we obtain perfect energy reflection. Thus remarkably, the stringy interface is completely tunable. As discussed in the End Matter, the averaged null energy condition (ANEC) is not violated as the additional conformal transformations in the quantum maps do not affect the total energy fluxes.
@@
Conclusions:- In this letter, we have demonstrated that the gravitational two-way junction including its stringy degrees of freedom can be decoded in terms of quantum maps. In absence of non-trivial stringy modes, these maps reduce to frequency independent universal maps [20, 31, 32, 21] that are realized by a defect operator, while the stringy modes imply that these maps should be generalized by composing them with appropriate conformal transformations. The full quantum maps preserve the conformal boundary condition on the interface. Particularly, the energy scattering involves independent and equal transmissions/reflections from both sides composed with an energy redistribution among the incoming/outgoing fluxes realized by conformal transformations. We have demonstrated the universality of these quantum maps by showing that they work for arbitrary input fluxes and an arbitrary background state (instead of the vacuum) on both sides of the interface within the universal sector.
Our calculations up to the linear order in the fluxes reveal that the holographic interface operator is a modification of (1) taking the form
$(L_{n,+}^{L}-L_{-n,-}^{L})I_{L,R}=I_{L,R}(\tilde{L}_{n,+}^{R}-\tilde{L}_{-n,-}^{R}),$ (38)
where $\tilde{L}_{\pm,m}$ are automorphisms of $L_{\pm,m}$ which preserve the Virasoro algebra. The understanding of the full non-linear gravitational problem, involving mixing of modes with different frequencies would be necessary to check if interface operators of the form (38) describe dual holographic junctions with stringy excitations.
@@
It would be also of profound interest to understand how explicit reconstruction of sub-regions of the gravitational junction from interface CFT sub-regions can work by generalizing the reformulation of bulk reconstruction in holography in terms of recovery maps of quantum error correcting codes [4, 7]. As a first step, it would be of interest to understand how stringy degrees of freedom of gravitational junctions influence entanglement in the dual CFT following [33, 34, 35].
@@
It would also be interesting to investigate the quantum null energy condition (QNEC) [36, 37, 38] in the holographic interfaces dual to gravitational junctions. The QNEC can impose non-trivial quantum thermodynamic bounds even if matter localized on bulk hypersurfaces satisfies the classical null energy condition [39, 40, 41] (see also [42] for a similar study in CFTs). The understanding of the link between quantum thermodynamic bounds and energy conditions is fundamental for the realization of quantum devices using interfaces in many-body systems.
@@ Acknowledgements.
Acknowledgments:- We thank Costas Bachas, Avik Banerjee and Giuseppe Policastro for valuable discussions. We especially thank Giuseppe Policastro for discussions related to the ANEC bounds and quantum maps. AC, AM and MM acknowledge support from FONDECYT postdoctoral grant no. 3230222, FONDECYT regular grant no. 1240955 and “Doctorado Nacional” grant no. 21250596 of La Agencia Nacional de Investigación y Desarrollo (ANID), Chile, respectively. AM gratefully acknowledges the hospitality of LPENS, where a substantial part of this work was carried out during his tenure as a CNRS invited professor.
@@
Checking achronal ANEC
Finally, the achronal ANEC is always satisfied for the generalized quantum maps at the linear order as the linearized conformal transformations do not change the energy fluxes if $a_{n}(\omega)$ vanishes as $\omega\rightarrow 0$ (see (32) as for instance). Under the following transformation
$g(x^{\pm})=x^{\pm}+\epsilon\mathbb{g}(x^{\pm}),$ (48)
the averaged null energy
$\int{\rm d}x^{\pm}T_{\pm\pm}$ (49)
is invariant at $\mathcal{O}(\epsilon)$ if
$\int{\rm d}x^{\pm}\mathbb{g}^{\prime\prime\prime}(x^{+})=0,$ (50)
when $T_{\pm\pm}$ are also $\mathcal{O}(\epsilon)$. For $\mathbb{t}_{d}$ and other variables to be well defined, we actually require that $a_{n}(\omega)$ vanishes at least as $\omega^{3}$ as $\omega\rightarrow 0$.